iq scores with a mean of 100.1 and a standard deviation of 16.9 fins the iq acore that seperates the bottom 25% from the too 75%
Answer #1

Given l=100.1 7:16.9 Separating bottom 95% from top 756 Area 0.25 Corresponds to 22-0.674 - A=0.5 / X = uto2 Aroon5 = 100.1+